# Indicating Robot Vision Capabilities with Augmented Reality

### v1

@@ 9 Conclusion
  • Funding: This work is supported by the Zhao Han’s startup fund provided by the Bellini College of Artificial Intelligence, Cybersecurity and Computing of the University of South Florida.
@@ A.2 Confidence Across Conditions
Figures 14-17 illustrate confidence distributions across different conditions. Figure 14 shows confidence ratings comparison among within-subject conditions (Baseline, Eye Socket, Near-Eye Blocks). Figure 15 adds the Extended Blocks condition, compare it with the three within-subject conditions. Figure 16 adds Blocks at Task condition, compare it with the three within-subject conditions. Figure 17 compares Extended Blocks and Blocks at Task.
@@ A.3 Cognitive Effort
Figures 18-21 illustrate workload distributions across different conditions. Figure 18 shows workload comparison among within-subject conditions (Baseline, Eye Socket, Near-Eye Blocks). Figure 19 adds the Extended Blocks condition, compare it with the three within-subject conditions. Figure 20 adds Blocks at Task condition, compare it with the three within-subject conditions. Figure 21 compares Extended Blocks and Blocks at Task.

### v2

@@ 9 Conclusion
  • Funding: This work is supported by Zhao Han’s startup fund provided by the Bellini College of Artificial Intelligence, Cybersecurity and Computing of the University of South Florida.
@@ A.2 Confidence Across Conditions
Figures 16-19 illustrate confidence distributions across different conditions. Figure 16 shows confidence ratings comparison among within-subject conditions (Baseline, Eye Socket, Near-Eye Blocks). Figure 17 adds the Extended Blocks condition, compare it with the three within-subject conditions. Figure 18 adds Blocks at Task condition, compare it with the three within-subject conditions. Figure 19 compares Extended Blocks and Blocks at Task.
@@ A.3 Cognitive Effort
Figures 20-23 illustrate workload distributions across different conditions. Figure 20 shows workload comparison among within-subject conditions (Baseline, Eye Socket, Near-Eye Blocks). Figure 21 adds the Extended Blocks condition, compare it with the three within-subject conditions. Figure 22 adds Blocks at Task condition, compare it with the three within-subject conditions. Figure 23 compares Extended Blocks and Blocks at Task.